# Neural stochastic Volterra equations: learning path-dependent dynamics

### v1

@@ Abstract.
Stochastic Volterra equations (SVEs) serve as mathematical models for the time evolutions of random systems with memory effects and irregular behaviour. We introduce neural stochastic Volterra equations as a physics-inspired architecture, generalizing the class of neural stochastic differential equations, and provide some theoretical foundation. Numerical experiments on various SVEs, like the disturbed pendulum equation, the generalized Ornstein–Uhlenbeck process and the rough Heston model are presented, comparing the performance of neural SVEs, neural SDEs and Deep Operator Networks (DeepONets).
@@ 1. Introduction
In the present work, we introduce neural stochastic Volterra equations as stochastic Volterra equations with coefficients parameterized by neural networks. They constitute a natural generalization of neural SDEs with the advantage that they are capable to represent time series with temporal dependency structures, which overcomes a limitation faced by neural SDEs. Hence, neural SVEs are suitable to serve as generative models for random dynamics with memory effects and irregular behaviour, even more irregular than neural SDEs. As theoretical justification for the universality of neural SVEs, we provide a stability result for general SVEs in Proposition 2.3, which can be combined with classical universal approximation theorems for neural networks [Cyb89, Hor91].
@@
We numerically investigate the supervised learning problem for prototypical Volterra type dynamics such as the disturbed pendulum equation [Øks03], the rough Heston model [EER19] and the generalized Ornstein–Uhlenbeck process [Vas12]. The performance of the neural SVE based models is compared to Deep Operator Networks (DeepONets) and to neural SDEs. Recall DeepONets are a popular class of neural learning algorithms for general operators on function spaces that were introduced in [LJP+21]. For the training process of the neural SVE we choose the Adam algorithm, as introduced in [KB14], which is known to be a well-suited stochastic gradient descent method for stochastic optimization problem.
The numerical study in Section 3 demonstrates that the presented neural SVE based methods outperform DeepONets significantly, see Table 1-Table 4. Especially, neural SVE based methods generalize much more efficiently as one can see at the good performance on the test sets, where neural SVEs are up to $20$ times more accurate than DeepONets. An advantage of neural SVEs are not only the physics-informed architecture, they are also time-resolution invariant, meaning that they can be trained and evaluated on arbitrary, possibly different time grid discretizations. Moreover, neural SVEs outperform neural SDE based models for random dynamics with a dependency structure, cf. Subsection 3.4.
Organization of the paper: In Section 2 we introduce neural stochastic Volterra equations and their theoretical background. The numerical experiments are presented in Section 3.
@@ 2. Neural stochastic Volterra equations
To define the notion of a (strong) $L^{p}$-solution, let $L^{p}(\Omega\times[0,T])$ be the space of all real-valued, $p$-integrable functions on $\Omega\times[0,T]$. We call an $(\mathcal{F}_{t})_{t\in[0,T]}$-progressively measurable stochastic process $(X_{t})_{t\in[0,T]}$ in $L^{p}(\Omega\times[0,T])$ on the given probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$, a (strong) $L^{p}$-solution of the SVE (2.1) if $\int_{0}^{t}(|K_{\mu}(t-s)\mu(s,X_{s})|+|K_{\sigma}(t-s)\sigma(s,X_{s})|^{2})\,\mathrm{d}s<\infty$ for all $t\in[0,T]$ and the integral equation (2.1) hold $\mathbb{P}$-almost surely. As usual, a strong $L^{1}$-solution $(X_{t})_{t\in[0,T]}$ of the SVE (2.1) is often just called solution of the SVE (2.1).
@@ 2.1. Neural SVEs
be seven feedforward neural networks (see [YYK15, Section 3.6.1]) that are parameterized by some common parameter $\theta$. Note that $L_{\theta}$ lifts the given initial value to the latent space $\mathbb{R}^{d_{h}}$, $\Pi_{\theta}$ is the readout back from the latent space to the space $\mathbb{R}^{d}$, and the other networks try to imitate their respectives in equation (2.1), on the latent $d_{h}$-dimensional space.
@@
The objective is to optimize $\theta$ as good as possible such that the generated paths are as close as possible to the given training paths. Therefore, one needs to solve a stochastic optimization problem at each training step. One typically chosen and well-suited stochastic gradient descent method for stochastic optimization problem is the Adam algorithm, introduced in [KB14]. The Adam algorithm is known to be computationally efficient, requires little memory, is invariant to diagonal rescaling of gradients and is well-suited for high-dimensional problems with regard to data/parameters.
@@ 2.3. Stability for SVEs
The mathematical reason that neural stochastic Volterra equations provide a suitable structure for learning the dynamics of general SVEs is the universal approximation property of neural networks, see e.g. [Cyb89, Hor91], and the stability result for SVEs, presented in this subsetions. More precisely, our stability result yields that if we approximate the kernels and coefficients of an SVE sufficiently well, we get a good approximation of the solution by the respective approximating solutions. To formulate the stability result, we need the following definitions and assumptions.
@@
$\|f\|_{\infty}:=\sup_{t\in[0,T],x\in\mathbb{R}^{d}}|f(t,x)|\quad\text{and}\quad\|g\|_{\infty}:=\sup_{t\in[0,T]}|g(x)|.$
As comparison to the SVE (2.1), we consider the SVE
@@
where $\tilde{g}\colon[0,T]\to\mathbb{R}$ is a continuous function, the coefficients $\tilde{\mu}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tilde{\sigma}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$, and the convolutional kernels $\tilde{K}_{\mu},\tilde{K}_{\sigma}\colon[0,T]\to\mathbb{R}$ are measurable functions.
@@
Suppose that $\|K_{\mu}\|_{q}+\|\tilde{K}_{\mu}\|_{q}<\infty$, $\|K_{\sigma}\|_{2\tilde{q}}+\|\tilde{K}_{\sigma}\|_{2\tilde{q}}<\infty$, $\|\mu\|_{\infty}+\|\tilde{\mu}\|_{\infty}+\|\sigma\|_{\infty}+\|\tilde{\sigma}\|_{\infty}<\infty$ and $\|g\|_{\infty}+\|\tilde{g}\|_{\infty}<\infty$.
@@
Let $\mu,\sigma,\tilde{\mu},\tilde{\sigma}\colon[0,T]\times\mathbb{R}\to\mathbb{R}$ be measurable functions such that:
@@
    $|\mu(t,x)-\mu(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C_{\mu,\sigma}|x-y|$
    holds for all $t\in[0,T]$ and $x,y\in\mathbb{R}$.
Based on these assumptions, we obtain the following stability result.
@@
(2.5) $\sup_{t\in[0,T]}\mathbb{E}[|X_{t}-\tilde{X}_{t}|^{p}]\leq C\Big{(}\|g-\tilde{g}\|_{\infty}^{p}+\|\mu-\tilde{\mu}\|_{\infty}^{p}+\|\sigma-\tilde{\sigma}\|_{\infty}^{p}+\|K_{\mu}-\tilde{K}_{\mu}\|_{2\tilde{q}}^{p}+\|K_{\sigma}-\tilde{K}_{\sigma}\|_{2\tilde{q}}^{p}\Big{)}.$
@@
$\leq C\bigg{(}\|g-\tilde{g}\|_{\infty}^{p}+|K_{\mu}-\tilde{K}_{\mu}|^{p}\int_{0}^{t}\big{(}1+\mathbb{E}\big{[}\big{|}X_{s}\big{|}^{p}\big{]}\big{)}\,\mathrm{d}s$
Using the regularity assumptions on $\mu$ and $\sigma$ (Assumption 2.2) and the boundedness of all moments of Volterra processes (see [PS23, Lemma 3.4]), we get
$\mathbb{E}\big{[}|X_{t}-\tilde{X}_{t}|^{p}\big{]}\leq C$ $\bigg{(}\|g-\tilde{g}\|_{\infty}^{p}+\|\mu-\tilde{\mu}\|_{\infty}^{p}+\|\sigma-\tilde{\sigma}\|_{\infty}^{p}+\|K_{\mu}-\tilde{K}_{\mu}\|_{2\tilde{q}}^{p}+\|K_{\sigma}-\tilde{K}_{\sigma}\|_{2\tilde{q}}^{p}\bigg{)}$
Applying Grönwall’s lemma leads to
$\mathbb{E}\big{[}|X_{t}-\tilde{X}_{t}|^{p}\big{]}\leq C$ $\bigg{(}\|g-\tilde{g}\|_{\infty}^{p}+\|\mu-\tilde{\mu}\|_{\infty}^{p}+\|\sigma-\tilde{\sigma}\|_{\infty}^{p}+\|K_{\mu}-\tilde{K}_{\mu}\|_{2\tilde{q}}^{p}+\|K_{\sigma}-\tilde{K}_{\sigma}\|_{2\tilde{q}}^{p}\bigg{)},$
which implies (2.5) by taking the supremum on the left-hand side. ∎
@@ 3. Numerical experiments
In this section we numerically investigate the supervised learning problem utilizing neural stochastic Volterra equations with the target to learn Volterra type dynamics such as the disturbed pendulum equation, the generalized Ornstein–Uhlenbeck process and the rough Heston model. The performance is compared to Deep Operator Networks and neural stochastic differential equations. For all the neural SVEs, we chose the latent dimensions $d_{h}=d_{K}=12$ which experimentally proved to be well-suited. We consider the interval $[0,T]$ for $T=5$ and discretize it equally-sized using the grid size $\Delta t=0.1$.
@@
Note that one major advantage of neural SVEs is that they are discretization invariant, i.e., that in the training procedure and also when evaluating a trained model, it does not matter if the input functions (realizations of the Brownian motion) are discretized by the same grid. In contrast, the DeepONet needs the same discretization for all functions in the training and in the evaluation data.
@@
Note that since DeepONet is not able to deal with random initial conditions, we use deterministic initial conditions $\xi=2$ in the DeepONet experiments. For neural SVEs, we use initial conditions $\xi\sim\mathcal{N}(2,0.2)$.
@@
Another advantage is that neural SVEs are time-resolution invariant, meaning that they can be trained and evaluated on arbitrary, possibly different time grid discretizations which is not possible for DeepONet.
@@ 3.1. Disturbed pendulum equation
where $\dot{B}_{t}=\frac{dB_{t}}{dt}$ is White noise for some standard Brownian motion $(B_{t})_{t\in[0,T]}$. Using the deterministic and the stochastic Fubini theorem, this system can be rewrite as stochastic Volterra equations
@@
$X(t)=x_{0}+tX^{\prime}(0)+\int_{0}^{t}(t-s)\frac{F(X(s))}{m}\,\mathrm{d}s+\int_{0}^{t}(t-s)\frac{\epsilon X_{s}}{m}\,\mathrm{d}B_{s}.$
@@
(3.1) $y_{t}=\xi+\int_{0}^{t}(t-s)y_{s}\,\mathrm{d}s+\int_{0}^{t}(t-s)y_{s}\,\mathrm{d}B_{s},\qquad t\in[0,T],$

### v2

@@ Abstract.
Stochastic Volterra equations (SVEs) serve as mathematical models for the time evolutions of random systems with memory effects and irregular behaviour. We introduce neural stochastic Volterra equations as a physics-inspired architecture, generalizing the class of neural stochastic differential equations, and provide some theoretical foundation. Numerical experiments on various SVEs, like the disturbed pendulum equation, the generalized Ornstein–Uhlenbeck process, the rough Heston model and a monetary reserve dynamics, are presented, comparing the performance of neural SVEs, neural SDEs and Deep Operator Networks (DeepONets).
@@ 1. Introduction
In the present work, we introduce neural stochastic Volterra equations as stochastic Volterra equations with coefficients parameterized by neural networks. They constitute a natural generalization of neural SDEs with the advantage that they are capable to represent time series with temporal dependency structures, which overcomes a limitation faced by neural SDEs. Hence, neural SVEs are suitable to serve as generative models for random dynamics with memory effects and irregular behaviour, even more irregular than neural SDEs. As theoretical justification for the universality of neural SVEs, we provide a stability result for general SVEs in Proposition 2.6, which can be combined with classical universal approximation theorems for neural networks [Cyb89, Hor91, KMFL20, KPT25]; cf. Remark 2.7.
@@
We numerically investigate the supervised learning problem for prototypical Volterra type dynamics such as the disturbed pendulum equation [Øks03], the rough Heston model [EER19], the generalized Ornstein–Uhlenbeck process [Vas12] and a model for the dynamics of monetary reserves [CFMS18]. The performance of the neural SVE based models is compared to Deep Operator Networks (DeepONets) and to neural SDEs. Recall DeepONets are a popular class of neural learning algorithms for general operators on function spaces that were introduced in [LJP+21]. For the training process of the neural SVE we choose the Adam algorithm, as introduced in [KB14], which is known to be a well-suited stochastic gradient descent method for stochastic optimization problem.
The numerical study in Section 3 demonstrates that the presented neural SVE based methods significantly outperform DeepONets; see Table 1-Table 3. In particular, neural SVE based methods generalize much more effectively, as evidenced by their strong performance on the test sets – neural SVEs are up to $20$ times more accurate than DeepONets. Moreover, neural SVEs also outperform neural SDE based models for random dynamics with dependency structures; cf. Subsection 3.5. These observations highlight the advantages of the physics-informed architecture of neural SVEs for supervised learning problems involving random systems with Volterra-type dynamics.
Organization of the paper: In Section 2 we introduce neural stochastic Volterra equations and their theoretical background. The numerical experiments are presented in Section 3. In Appendix A we present the postponed proofs regarding the stability of stochastic Volterra equations.
@@ 2. Neural stochastic Volterra equations
To define the notion of a (strong) $L^{p}$-solution, let $L^{p}(\Omega\times[0,T])$ be the space of all real-valued, $p$-integrable functions on $\Omega\times[0,T]$. We call an $(\mathcal{F}_{t})_{t\in[0,T]}$-progressively measurable stochastic process $(X_{t})_{t\in[0,T]}$ in $L^{p}(\Omega\times[0,T])$, on the given probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$, a (strong) $L^{p}$-solution of the SVE (2.1) if $\int_{0}^{t}(|K_{\mu}(t-s)\mu(s,X_{s})|+|K_{\sigma}(t-s)\sigma(s,X_{s})|^{2})\,\mathrm{d}s<\infty$ for all $t\in[0,T]$ and the integral equation (2.1) hold $\mathbb{P}$-almost surely. As usual, a strong $L^{1}$-solution $(X_{t})_{t\in[0,T]}$ of the SVE (2.1) is often just called solution of the SVE (2.1).
@@ 2.1. Neural SVEs
be seven feedforward neural networks (see [YYK15, Section 3.6.1]) that are parameterized by some common parameter $\theta$. Note that $L_{\theta}$ lifts the given initial value to the latent space $\mathbb{R}^{d_{h}}$, $\Pi_{\theta}$ is the readout back from the latent space to the space $\mathbb{R}^{d}$, and the other networks try to imitate their respectives in Equation (2.1) on the latent $d_{h}$-dimensional space.
@@
The objective is to optimize $\theta$ as good as possible such that the generated paths are as close as possible to the given training paths. Therefore, one needs to solve a stochastic optimization problem at each training step. One typically chosen and well-suited stochastic gradient descent method for stochastic optimization problems is the Adam algorithm, introduced in [KB14]. The Adam algorithm is known to be computationally efficient, requires little memory, is invariant to diagonal rescaling of gradients and is well-suited for high-dimensional problems with regard to data/parameters.
@@ 2.3. Stability for SVEs
The mathematical reason that neural stochastic Volterra equations provide a suitable structure for learning the dynamics of general SVEs is the universal approximation property of neural networks, see e.g. [Cyb89, Hor91, KL20, KPT25], and the stability result for SVEs, presented in this subsection. More precisely, our stability result yields that if we approximate the kernels and coefficients of an SVE sufficiently well, we get a good approximation of the solution by the respective approximating solutions. To formulate the stability result, we need the following definitions and assumptions.
@@
$\|f\|_{\infty}:=\sup_{t\in[0,T],x\in\mathbb{R}^{d}}|f(t,x)|\quad\text{and}\quad\|g\|_{\infty}:=\sup_{t\in[0,T]}|g(t)|.$
As approximation of the SVE (2.1), we consider a sequence of SVEs, given by
(2.3) $X^{n}_{t}=\xi g_{n}(t)+\int_{0}^{t}K_{\mu,n}(t-s)\mu_{n}(s,X^{n}_{s})\,\mathrm{d}s+\int_{0}^{t}K_{\sigma,n}(t-s)\sigma_{n}(s,X^{n}_{s})\,\mathrm{d}B_{s},$
for $t\in[0,T]$ and $n\in\mathbb{N}$. We make the following assumptions on the kernels $K_{\mu,n},K_{\sigma,n}$, the coefficients $\mu_{n},\sigma_{n}$ and the initial conditions $g_{n}$.
Assumption 2.1.
The initial conditions $g,g_{n}\colon[0,T]\to\mathbb{R}^{d}$ and kernels $K_{\mu},K_{\sigma}\colon[0,T]\to\mathbb{R}$ and $K_{\mu,n},K_{\sigma,n}\colon[0,T]\to\mathbb{R}$ for $n\in\mathbb{N}$ satisfy the following conditions: There are constants $\gamma\in(0,\frac{1}{2}]$, $\varepsilon>0$ and $L>0$, such that
  1. (i)
    for all $n\in\mathbb{N}$ the measurable functions $K_{\mu,n},K_{\sigma,n}\colon[0,T]\to[0,\infty)$ fulfill
    $\int_{0}^{T-h}|K_{\mu,n}(h+r)-K_{\mu,n}(r)|^{1+\varepsilon}\,\mathrm{d}r+\int_{0}^{h}|K_{\mu,n}(r)|^{1+\varepsilon}\,\mathrm{d}r\leq Lh^{\gamma(1+\varepsilon)},$
    $\int_{0}^{T-h}|K_{\sigma,n}(h+r)-K_{\sigma,n}(r)|^{2+\varepsilon}\,\mathrm{d}r+\int_{0}^{h}|K_{\sigma,n}(r)|^{2+\varepsilon}\,\mathrm{d}r\leq Lh^{\gamma(2+\varepsilon)},$
    for all $h\in[0,T]$;
  2. (ii)
    it holds that
    $\int_{0}^{T}|K_{b,n}(s)-K_{b}(s)|\,\mathrm{d}s\to 0\text{ as }n\to\infty$
    and
    $\int_{0}^{T}|K_{\sigma,n}(s)-K_{\sigma}(s)|^{2+\varepsilon}\,\mathrm{d}s\to 0\text{ as }n\to\infty;$
  3. (iii)
    $g$ and $g_{n}$ are $\gamma$-Hölder-continuous.
Assumption 2.2.
Let $\mu,\mu_{n}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma,\sigma_{n}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$, $n\in\mathbb{N}$, be measurable functions such that:
  1. (i)
    $\mu,\sigma$ and $\mu_{n},\sigma_{n}$ are (uniformly) of linear growth, i.e. there is a constant $C_{\mu,\sigma}>0$ such that
    $\sup_{n\in\mathbb{N}}|\mu_{n}(t,x)|+|\sigma_{n}(t,x)|+|\mu(t,x)|+|\sigma(t,x)|\leq C_{\mu,\sigma}(1+|x|),$
    for all $t\in[0,T]$ and $x\in\mathbb{R}^{d}$.
  2. (ii)
    For any compact subset $K\subset\mathbb{R}^{d}$ we have
    $\lim_{n\to\infty}\sup_{t\in[0,T]}\sup_{x\in K}|\sigma_{n}(t,x)-\sigma(t,x)|+|\mu_{n}(t,x)-\mu(t,x)|+|g_{n}(t)-g(t)|=0.$
Based on the aforementioned assumptions, we obtain the following stability result for stochastic Volterra equations, generalizing the classical stability result for ordinary stochastic differential equations proven in [KN88].
Theorem 2.3.
Suppose Assumption 2.1 and Assumption 2.2 and let $\xi\in L^{p}(\Omega)$ with $p>\max\{\frac{1}{\gamma},\frac{4+2\varepsilon}{\varepsilon}\}$. Moreover, suppose that there are unique $L^{p}$-solutions $(X_{t})_{t\in[0,T]}$ and $(X^{n}_{t})_{t\in[0,T]}$ to the SVEs (2.1) and (2.3), for $n\in\mathbb{N}$, respectively. Then, one has
$\lim_{n\to\infty}\mathbb{E}\bigg{[}\sup_{t\in[0,T]}|X^{n}_{t}-X_{t}|^{2}\bigg{]}=0.$
Proof.
See Appendix A. ∎
Note that the assumption on the existence of unique solutions can be ensured by postulating the coefficients to be Lipschitz continuous, see e.g. [Wan08]. However, for instance, in a one-dimensional setting a unique solution can also be obtained for Hölder continuous diffusion coefficients, see e.g. [AJEE19, PS23b].
Assuming that the coefficients of a SVE are Lipschitz continuous, one can quantify the stability result of Theorem 2.3, as we shall present below. To that end, we consider, as comparison to the SVE (2.1), the SVE
@@
where $\tilde{g}\colon[0,T]\to\mathbb{R}$ is a continuous function, and where the coefficients $\tilde{\mu}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\tilde{\sigma}\colon[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$, and the convolutional kernels $\tilde{K}_{\mu},\tilde{K}_{\sigma}\colon[0,T]\to\mathbb{R}$ are measurable functions.
@@
Suppose that $\|K_{\mu}\|_{q}+\|\tilde{K}_{\mu}\|_{q}<\infty$, $\|K_{\sigma}\|_{2\tilde{q}}+\|\tilde{K}_{\sigma}\|_{2\tilde{q}}<\infty$, and $\|g\|_{\infty}+\|\tilde{g}\|_{\infty}<\infty$.
@@
Let $\mu,\tilde{\mu}\colon[0,T]\times\mathbb{R}\to\mathbb{R}^{d}$ and $\sigma,\tilde{\sigma}\colon[0,T]\times\mathbb{R}\to\mathbb{R}^{d\times m}$ be measurable functions such that:
@@
    $|\tilde{\mu}(t,x)-\tilde{\mu}(t,y)|+|\tilde{\sigma}(t,x)-\tilde{\sigma}(t,y)|\leq C_{\mu,\sigma}|x-y|$
    holds for all $t\in[0,T]$ and $x,y\in\mathbb{R}^{d}$.
Based on these assumptions, we obtain the following stability result for stochastic Volterra equations with Lipschitz continuous coefficients. For related stability results in the context ordinary stochastic differential equations, we refer to [KPT25, Section 3] and the references therein.
@@
(2.6) $\sup_{t\in[0,T]}\mathbb{E}[|X_{t}-\tilde{X}_{t}|^{p}]\leq C\Big{(}\|g-\tilde{g}\|_{\infty}^{p}+\|\mu-\tilde{\mu}\|_{\infty}^{p}+\|\sigma-\tilde{\sigma}\|_{\infty}^{p}+\|K_{\mu}-\tilde{K}_{\mu}\|_{q}^{p}+\|K_{\sigma}-\tilde{K}_{\sigma}\|_{2\tilde{q}}^{p}\Big{)}.$
@@
See Appendix A. ∎
Remark 2.7.
The stability results, presented in Theorem 2.3 and Proposition 2.6, demonstrate the universality of neural stochastic Volterra equations like (2.1). Indeed, the unique solution of any general stochastic Volterra equations can be approximated arbitrary well by solutions of neural stochastic Volterra equations, assuming that the associated neural network converges in a suitable sense. For the coefficients the required suitable convergence is the local uniform convergence subject to a uniform global linear growth constraint. As shown in [KPT25], many frequently used classes of neural networks do allow for such convergence, in particular, neural networks based on LipSwish activation functions, as we use in the numerical experiments below. For the kernels, the suitable type of convergence is an $L^{p}$-convergence on the compact interval $[0,T]$. $L^{p}$-type universal approximation theorem can already be found in the classical work of [Hor91]; we also refer to [KL20] and the references therein.
@@ 3. Numerical experiments
In this section, we numerically investigate the supervised learning problem utilizing neural stochastic Volterra equations aiming to learn Volterra type dynamics such as the disturbed pendulum equation, the generalized Ornstein–Uhlenbeck process, a model for the dynamics of monetary reserves, and the rough Heston model. The performance is compared to Deep Operator Networks and neural stochastic differential equations. For all the neural SVEs, we chose the latent dimensions $d_{h}=d_{K}=12$ which experimentally proved to be well-suited. We consider the interval $[0,T]$ for $T=5$ and discretize it equally-sized using the grid size $\Delta t=0.1$.
@@
Note that since DeepONet is not able to deal with random initial conditions, we use deterministic initial conditions $\xi=2$ in the DeepONet experiments. For neural SVEs we use initial conditions $\xi\sim\mathcal{N}(2,0.2)$ unless stated otherwise.
@@ 3.1. Disturbed pendulum equation
where $\dot{B}_{t}=\frac{dB_{t}}{dt}$ is White noise for some standard Brownian motion $(B_{t})_{t\in[0,T]}$. Using the deterministic and the stochastic Fubini theorem, this system can be rewritten as stochastic Volterra equations
@@
$X(t)=x_{0}+tX^{\prime}(0)+\int_{0}^{t}(t-s)\frac{F(X(s))}{m}\,\mathrm{d}s+\int_{0}^{t}(t-s)\frac{\varepsilon X_{s}}{m}\,\mathrm{d}B_{s}.$
@@
(3.1) $y_{t}=\xi-\int_{0}^{t}(t-s)y_{s}\,\mathrm{d}s+\int_{0}^{t}(t-s)y_{s}\,\mathrm{d}B_{s},\qquad t\in[0,T],$
@@
To highlight the performance of Neural SVE in a more complex setting, we take a nonlinear coefficient $\mu$ and consider
(3.2) $y_{t}=\xi-\int_{0}^{t}(t-s)\sin(y_{s})\,\mathrm{d}s+\int_{0}^{t}0.4(t-s)y_{s}\,\mathrm{d}B_{s},\qquad t\in[0,T].$
The results are presented in Table 2. It stands out immediately that for $n=100$ the performance in the test set is far worse than in the training set. Most solutions remain in the range of $[-\pi,\pi]$. Yet, in some rare cases, solutions may explode outside this range. An example of this can be seen in Figure 3.1. In the training set this happens too rarely for the neural network to learn the functions outside of $[-\pi,\pi]$ precisely. Table 2 clearly shows that this discrepancy disappears when the training set is sufficiently large. In the case of $n=2000$ one also sees that the neural network can learn these complex functions as good as in the prior example. Then the plots for the test set look similar to the training set plots of the Neural SVE trained with $n=100$ trajectories.
Neural SVE Train set Test set
$n=100$ $0.007$ $0.028$
$n=500$ $0.012$ $0.014$
$n=2000$ $0.007$ $0.006$
DeepONet Train set Test set
$n=100$ $0.004$ $0.23$
$n=500$ $0.006$ $0.13$
$n=2000$ $0.004$ $0.06$
Table 2. Mean relative $L^{2}$-losses after training for the nonlinear disturbed pendulum equation (3.2).
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set DeepONet:
Training set DeepONet:
Test set
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2. Sample neural SVE and DeepONet paths from the training and the test set for the disturbed pendulum equation with nonlinear drift and $n=100$. Blue are the original paths and orange the learned approximations.
3.2. Rough Heston equation
Neural SVE Train set Test set
$n=100$ $0.003$ $0.003$
$n=500$ $0.0025$ $0.0028$
$n=2000$ $0.0015$ $0.0017$
DeepONet Train set Test set
$n=100$ $0.035$ $0.13$
$n=500$ $0.004$ $0.037$
$n=2000$ $0.003$ $0.014$
Table 3. Mean relative $L^{2}$-losses after training for the rough Heston equation (3.3).
{tblr}
—m3cm—m3cm—m3cm—m3cm— Neural SVE:
Training set Neural SVE:
Test set DeepONet:
Training set DeepONet:
Test set
Refer to caption Refer to caption Refer to caption Refer to caption